OpenStudy (sanic):

The seventh-grade class is building target areas for a PE activity. The bases for the game will be in a circular shape. The diameter of each circle is 3 feet. Approximately, how many square feet of the turf need to be painted for a base circle? Use 3.14 for π and round your answer to the nearest tenth. 7.1 square feet 9.4 square feet 18.8 square feet 28.3 square feet Will fan and medal

OpenStudy (welshfella):

area of each circular base = pi r^2 = 3.14 * 3^2

OpenStudy (sanic):

ok

OpenStudy (sanic):

@welshfella

OpenStudy (sanic):

would it be 9.4

OpenStudy (welshfella):

oh yes Sorry the diameter is 3 feet so the radius is 1.5 feet so area = 3.14 * 1.5^2 I misread the question

OpenStudy (sanic):

so what i got is not the answer

OpenStudy (welshfella):

3.14 * 1.5^2 = 3.14 * 2.25 = 7.1 sq ft

OpenStudy (sanic):

ok now i see it thank u

OpenStudy (welshfella):

yw